OpenStudy (anonymous):

Factor completely: 9x^2 + 42x + 49 (3x + 7)(3x − 7) (9x − 7)(9x − 7) (3x + 7)(3x + 7) (9x + 7)(9x + 7)

OpenStudy (anonymous):

Its c right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\(\large 9x^2 + 42x + 49=(3x)^2+2\times 3x\times7+7^2=(3x+7)^2\)

OpenStudy (anonymous):

Thanks!